How many cubed pieces of fudge that are 3 inches on an edge can be packed into a Christmas tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to be closed?
On a Map, 1 inch represents 20 miles. The distance between 2 towns is 6 1/5 inches. How many miles are actually between the two towns.
These two problems are really kicking my butt, please help your my last resort, please anyone.

Answer by Fombitz(32388)

1. How many cubed pieces of fudge that are 3 inches on an edge can be packed into a Christmas tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to be closed?
Look for the closest multiple of 3 that does not go over your measure for depth, width, and height.
9 inches deep - 3 pieces x 3" deep
12 inches wide - 4 pieces x 3" wide
8 inches high - 2 pieces x 3" high (6 inches, with 2 inches of air left).
For the total number of pieces, multiply the number of pieces depp, wide, and high.
Total number of pieces = 3x4x2 = 24 pieces of fudge can fit in the tin.
2. On a Map, 1 inch represents 20 miles. The distance between 2 towns is 6 1/5 inches. How many miles are actually between the two towns.
1"=20 miles
6(1")=6(20 miles)
6"=120 miles
1/5(1")=1/5(20 miles)
1/5"=4 miles
6"+1/5"=120+4=124 miles
6 1/5"=124 miles.